john is 4 years less than twice his little brother zachs age. if the product of their ages is 126, how old is each brother?
Answer by stanbon(75887)

john is 4 years less than twice his little brother zachs age. if the product of their ages is 126, how old is each brother?
-------
Equations:
j = 2b-4
j*b = 126
------
Substitute for "j" and solve for "b":
(2b-4)b = 126
2b^2-4b-126 = 0
b^2 - b - 63 = 0
b = 8.45.. year (brother's age)
----
Solve for "j":
j = 2(8.45)-4
j = 12.9 years (john's age)
=============================
Cheers,
Stan H.
==============